OpenStudy (anonymous):

solve the following equation by first multiplying both sides by an appropriate power of 10 so that the equation contains intigers only. 5x-22.8=x solve the following equation by first multiplying both sides by an appropriate power of 10 so that the equation contains intigers only. 5x-22.8=x @Mathematics

hero (hero):

5x - x = 22.8 4x = 22.8 40x = 228 x = 228/40

OpenStudy (mertsj):

50x-228=10x

hero (hero):

This is useless because x still gives a decimal

OpenStudy (anonymous):

the answer is 5.7

OpenStudy (mertsj):

It supposedly makes the equation easier for the beginner because it contains only integers and also gives practice in multiplying by powers of ten. Whether it makes sense or not is academic. It is a requirement of the problem solution.

OpenStudy (anonymous):

i need the answer

hero (hero):

You just gave yourself the answer. Then you say you need the answer. I'm confused

OpenStudy (mertsj):

50x-228 = 10x Subtract 10x from both sides and add 228 to both sides: 40x = 228 Divide both sides by 40 x = 228/40 Either do the long division and get 5.7 or pick up a calculator and do 228 divided by 40. 228/40 means 228 divided by 40

jhonyy9 (jhonyy9):

because 228/40 = 5,7